OpenStudy (vortish):

factor the trinomial completely 4u^2+4u-15

hartnn (hartnn):

tried this ?

OpenStudy (vortish):

i do not think this can be factored i think it is a prime

OpenStudy (vortish):

i does not make sense 4 can not be factored in t o15

hartnn (hartnn):

how you could conclude that ? did you find out the value of determinant b^2-4ac ?

hartnn (hartnn):

actually you need to find 2 numbers with product 4*(-15)=60 and sum 4

hartnn (hartnn):

4*(-15)=-60

OpenStudy (vortish):

so what your saying i have to find a number that -60 will = 4

hartnn (hartnn):

2 numbers whose product is -60 and whose sum is 4 (here, one number is negative)

OpenStudy (vortish):

-1086=-60 10*-6=-60 10-6=4

hartnn (hartnn):

-1086 ?? 10 and -6 are correct, so split +4u into 10u-6u now try...

hartnn (hartnn):

ohh, 10 & 6

OpenStudy (vortish):

(u+10)(u-6)

hartnn (hartnn):

huh ? no... 4u^2 +10u -6u -15 =0 factor out 2u from 1st 2 terms and -3 from last 2 terms, what u get ?

OpenStudy (vortish):

2u+4u -5

hartnn (hartnn):

?? (4u^2 +10u) + (-6u -15) =0 2u (2u+5) -3 (2u+5) =0 got this ?

OpenStudy (vortish):

yes sir

hartnn (hartnn):

now factor out 2u+5 from that

OpenStudy (vortish):

(2u-3)=0

hartnn (hartnn):

2u+5 just disappeared :O :P (2u+5)(2u-3) is your factored form :)

OpenStudy (vortish):

so you can only factor out one of the (2u+5) out of the expression

OpenStudy (vortish):

so the final answer would be (2u+5)(2u-3)

hartnn (hartnn):

its like ab+ac = a(b+c) (2u+5)*2u + (2u+5)*(-3) = (2u+5)(2u-3) and yes.

OpenStudy (vortish):

ok thank you sir

hartnn (hartnn):

welcome ^_^